OpenStudy (anonymous):

12 + 0.35x = 20.05 A. 91.5 B. 57.3 C. 2.8175 D. 23

OpenStudy (anonymous):

D. 23

OpenStudy (anonymous):

12+0.35x=20.05 12-12+0.35x=20.05-12.00 0.35x=8.05 .35/.35x=8.05/.35 x=23

OpenStudy (anonymous):

Hope that helps :)